OpenStudy (anonymous):

The tenth term of an arithmetic sequence is (57/2), and the second term is (9/2). Find the first term.

OpenStudy (nikvist):

\[a_{10}=a_1+9d=28.5\]\[a_2=a_1+d=4.5\]\[a_1+9d-(a_1+d)=8d=24\quad\Rightarrow\quad d=3\]\[a_1=a_2-d=4.5-3=1.5\]